# This work was supported in part by Hong Kong Innovation and Technology Commission Funding Administrative System II (ITF Ref. No. GHP/110/20GD) and the National Natural Science Foundation of China (62192714, U21B2006).(Corresponding author: Hongwei Liu; Chang Gao.)

### v2

@@ 5.2 The detection performance analysis
To demonstrate the superiority of the proposed GLP-MFD algorithm, a comparison is conducted with the traditional search-before-detect MFD method (NDP) and the MFD method that combines model-driven track initiation with non-coherent integration (MTS-NCI). The NDP method constructs potential tracks using kinematic constraints derived from target position measurements and makes track detection through non-coherent integration of track energy followed by comparison with a second threshold. The MTS-NCI method leverages Doppler information, creating kinematic constraints based on the coupling relationship between position and Doppler measurements to obtain potential tracks. By incorporating Doppler information, the separability between targets and false alarms is enhanced. Subsequently, the method accumulates track energy incoherently and makes track detection based on a second threshold. It is important to note that the NDP method does not utilize Doppler information. In the experimental scenario set for this work, if the NDP method uses the primary threshold corresponding to the initial false alarm probability $P_{fa,1}=10^{-3}$, it will lead to a combinatorial explosion, resulting in an unmanageable computational load. Additionally, due to its poor separability between targets and false alarms, an excessive number of false alarms will adversely impact correct associations of target measurements, degrading the performance of target detection. Therefore, for the NDP method, we employ the primary threshold corresponding to the initial false alarm probability $P_{fa,1,NDP}=10^{-4}$. The proposed GLP-MFD method and the MTS-NCI method both utilize the primary threshold corresponding to $P_{fa,1}=10^{-3}$. To ensure a fair comparison, all methods guarantee that the final false alarm probability $P_{fa,2}$ have to be $10^{-6}$ after processing multi-frame data.
@@
Under the same SNR conditions, the relationship between target detection probabilities and sliding window numbers for aforementioned methods is shown in Fig. 5. $P_{D,CA-CFAR}$ represents the target detection probability in a single frame using the cell-averaging constant false alarm rate (CA-CFAR) algorithm under a false alarm probability of $10^{-6}$. $P_{D,UPB}$ denotes the upper limit of detection performance, defined as the probability that a candidate track exists where all plots are target observations across multiple frames. It can be observed from Fig. 5 that as the sliding window number increases, the detection probability of multi-frame detection methods gradually improves. In most cases, multi-frame detection methods achieve better performance compared to CA-CFAR. However, the detection performance of NDP is mostly lower than that of CA-CFAR when the number of sliding windows is small. On one hand, the NDP method does not utilize Doppler information, resulting in poor distinction between targets and false alarms, and the exclusion of false tracks mainly relies on energy accumulation. On the other hand, the primary threshold is set corresponding to $P_{fa,1,NDP}=10^{-4}$, which restricts the upper limit of detection performance. In contrast, the MTS-NCI method, which incorporates Doppler information, demonstrates better separability between targets and false alarms, outperforming NDP. However, weak targets can diminish the separability and setting $P_{fa,1}=10^{-3}$ leads to a large number of false alarms. Consequently, when the SNR is 7 dB, the detection performance with a single sliding window is slightly lower than that of CA-CFAR. Whereas, under high SNR conditions, its performance consistently surpasses CA-CFAR.
@@
The proposed GLP-MFD algorithm integrates multi-dimensional information of target observations and spatio-temporal coupling information across multiple frames, approaching a near-optimal detector for fusing features on different attributes and achieving the highest detection performance. Moreover, its superiority becomes more apparent at low SNR compared with NDP and MTS-NCI. It is important to note that in Monte Carlo experiments, situations may arise where a target is not detected by $P_{D,UPB}$, yet the proposed algorithm concludes that the target has been correctly detected. This is because GLP-MFD defines correct detection based on the OSPA distance, which can occasionally obtain the output track where false alarms are present and the number of target plots is less than $M$. According to the definition of $P_{D,UPB}$, this do not count as a correct detection. However, the compliance of the condition shown in (26) indicates that the false alarms in the track do not affect subsequent target tracking, leading the consideration of correct detection. In practical engineering applications, such high-quality false alarms can also be advantageous for target track initiation.
@@
To completely show the experimental scenario and the overall performance of GLP-MFD, the processing results for one exemplar run are shown in Fig. 6. The observation area of interest is from $100$km to $300$km and from $-60^{\circ}$ to $60^{\circ}$ containing a total of 16 targets. The detection result corresponding to the primary threshold is shown in 6(a), with red solid lines indicating the tracks with correct target states. The processing results of the proposed GLP-MFD is displayed in 6(b), where the red solid lines represent the target tracks output from the network, black solid points indicate the false tracks, and blue squares denote the position measurements of targets.
@@
In the processing result, 3 false tracks and 14 target tracks are initiated, while two target tracks are not detected. The retention of false tracks may be due to their alignment with the multi-frame correlation characteristics of targets. However, they can be further eliminated through false track discrimination methods. Among 14 initiated target tracks, 13 are identified as correct target tracks. The enlarged track in the figure has a large OSPA distance from the true target track due to the presence of false alarms, which affects the target track direction and may impact subsequent tracking. Therefore, it cannot be considered correctly detected. Nevertheless, these erroneous target tracks do not impose excessive negative effects. Over continuous multi-frame processing, these incorrect initiations will gradually be filtered out, while the correct target plots within the erroneous tracks will remain and appear in the newly initiated correct tracks. In summary, the proposed GLP-MFD algorithm demonstrates strong detection performance for weak targets and can accurately detect target tracks while effectively suppressing false tracks.
5.3 Interpretable analysis
@@
To explore how MFLPN performs integrated track search and track detection based on input features, we analyze the importance of each input feature on the network output. The input features include signal-to-noise ratio (SNR), Doppler channel (DC), chronological information (CI), range-Doppler map (RDM), spatio-temporal coupling (STC) and the differences in Doppler channels (DCD). To analyze the contribution of these features to the network, techniques such as feature occlusion and feature permutation can be employed to eliminate the dependency between samples and their corresponding features. And the decline in network performance can be observed to confirm the importance of each feature[46],[47],[48]. The core task of this work is a link prediction task, which determines whether the associations between measurements originate from targets. To clarify the network decision criteria under different conditions, we use the extent of accuracy decline in link predictions caused by eliminating sample-feature dependencies as a reference metric. The box plots are used to explain the feature importance for the network[49]. The results of association decision in high SNR targets, low SNR targets, and false alarms are shown in Fig. 7.
@@
The results indicate that for high SNR targets, SNR is the most significant feature influencing performance. This is because a higher SNR enhances the separability between targets and false alarms, making it a key characteristic for distinguishing them. The next important features are the spatio-temporal coupling and the differences in Doppler channels. Effectively utilizing these features can enhance the accuracy of track initiation for targets with complex physical kinematics. For low SNR targets, although SNR remains the most influential feature, its impact reduces compared to high SNR targets. To compensate, the network increasingly relies on the range-Doppler map, which reflects not only the intensity of echo signals but also the distribution of target echo energy across the range and Doppler dimensions. This structural information facilitates deeper feature extraction by the network. Furthermore, since the network accuracy in making association decisions for low SNR targets is lower than that for high SNR targets, the range of outliers tends to be larger. In the case of associations between false alarms, the information extracted from SNR of noise measurements is limited, prompting the network to rely more on the range-Doppler map, which contains rich structural information. Additionally, due to the randomness in the distribution of false alarms, the contribution of the differences in Doppler channels decreases, while the spatio-temporal coupling feature becomes critical for the network decision-making. Specifically, if the deviation in the coupling between measured Doppler and position information is small, the association is more likely to stem from a target. Conversely, it is more likely to originate from false alarms.
@@
Thus, through extensive data training, the proposed MFLPN can automatically adjust its association criteria based on data features under different conditions, thereby adaptively selecting corresponding features to make association decisions for observations from various sources. Compared to traditional model-driven methods, the proposed method effectively integrates multi-dimensional information on different attributes to achieve near-optimal unified track search and track detection in multiple frames.
@@ 6 SUMMARY
To address the challenges in traditional single-frame detection methods, which struggle to balance detection performance and system processing burden in weak target detection scenarios, and to overcome the limitations of traditional multi-frame detection methods, where sequential track search and detection are hindered by model mismatches and insufficient utilization of echo information, we propose a multi-frame detection algorithm via graph neural network-based link prediction. The proposed algorithm constructs observation association graphs through a relative low primary threshold, leveraging graph neural networks to fully integrate various information, such as Doppler information, echo structures, and spatio-temporal coupling of multi-frame observations, to complete track search and track detection in a unified manner.
Experiments show that, compared to traditional single-frame and multi-frame detection methods, the proposed algorithm effectively suppresses false tracks while enhancing target detection performance. Moreover, the interpretable analysis reveals that the designed network effectively integrates the input features, enabling precise differentiation of associations between targets and false alarms. However, the multi-frame detection process of the proposed algorithm still involves relatively high computational load. Even though the low threshold eliminates observations unlikely from targets compared to raw data and the integration process follows a relatively straightforward logic, the computational burden during multi-frame joint processing remains substantial. Future work may consider further enhancing computational efficiency without compromising weak target detection performance.

### v3

@@ 5.3 The detection performance analysis
To demonstrate the superiority of the proposed GLP-MFD, a comparison is conducted with the three-stage MFD method using the novel dynamic programming (NDP)[18], and the method that combines Doppler-aided track formation with non-coherent integration (DTF-NCI)[23]. The NDP method constructs potential tracks using kinematic constraints derived from target position measurements. And it makes track detection through non-coherent integration of track energy followed by a comparison with a final detection threshold. The DTF-NCI method designs kinematic constraints based on Doppler information to obtain potential tracks. By incorporating Doppler information, the separability between targets and false alarms is enhanced. Subsequently, the method accumulates track energy incoherently and makes track detection based on a final detection threshold. It is important to note that the NDP method does not utilize Doppler information. In the experimental scenario set for this work, if the NDP method uses the primary threshold corresponding to the initial false alarm probability $P_{fa,1}=10^{-3}$, it will lead to a combinatorial explosion, resulting in an unmanageable computational load. Additionally, due to its poor separability between targets and false alarms, an excessive number of false alarms will adversely impact correct associations of target measurements, degrading the performance of target detection. Therefore, for the NDP method, we employ the primary threshold corresponding to the initial false alarm probability $P_{fa,1,NDP}=10^{-4}$. The proposed GLP-MFD method and the DTF-NCI method both utilize the primary threshold corresponding to $P_{fa,1}=10^{-3}$. To ensure a fair comparison, all methods guarantee that the final false alarm probability $P_{fa,2}$ have to be $10^{-6}$ after processing multi-frame data.
Under the same SNR conditions, the relationship between target detection probabilities and sliding window numbers for aforementioned methods is shown in Fig. 8. $P_{D,CA-CFAR}$ represents the target detection probability in a single frame using the cell-averaging constant false alarm rate (CA-CFAR) method under a false alarm probability of $10^{-6}$. $P_{D,UPB}$ denotes the upper limit of detection performance, defined as the probability that a candidate track exists where all plots are target observations across multiple frames. It can be observed from Fig. 8 that as the sliding window number increases, the detection probability of multi-frame detection methods gradually improves. In most cases, multi-frame detection methods achieve better performance compared to CA-CFAR. However, the detection performance of NDP is mostly lower than that of CA-CFAR when the number of sliding windows is small. On one hand, the NDP method does not utilize Doppler information, resulting in poor distinction between targets and false alarms. In addition, the exclusion of false tracks mainly relies on energy accumulation. On the other hand, the primary threshold is set corresponding to $P_{fa,1,NDP}=10^{-4}$, which restricts the upper limit of detection performance. In contrast, the DTF-NCI method, which incorporates Doppler information, demonstrates better separability between targets and false alarms, outperforming NDP. However, weak targets and the dense false alarm environment can diminish the separability. Consequently, when the SNR is 7 dB, the detection performance of DTF-NCI with a single sliding window is slightly lower than that of CA-CFAR. Whereas, under high SNR conditions, its performance consistently surpasses CA-CFAR.
@@
The CA-CFAR method exhibits poor detection performance for weak targets because it only uses single-frame information. Although the NDP method utilizes multi-frame information for target detection, it merely uses position measurements when searching potential tracks and does not fully utilize Doppler information, which reduces the separability between targets and false alarms. The DTF-NCI method incorporates Doppler information to construct kinematic constraints when searching potential tracks, but such relatively strict physical constraints may lead to target missed associations, resulting in performance loss. Additionally, the non-coherent integration track detector does not fully exploit the multi-dimensional information on different attributes across frames, which leads to a performance bottleneck [58]. The proposed GLP-MFD algorithm integrates multi-dimensional information of target observations and spatio-temporal coupling information across multiple frames, approaching a detector for fusing features on different attributes and achieving the highest detection performance. Moreover, its superiority becomes more apparent at low SNR conditions compared with NDP and DTF-NCI. It is important to note that in Monte Carlo experiments, situations may arise where a target is not detected by $P_{D,UPB}$, yet the proposed algorithm concludes that the target has been correctly detected. This is because GLP-MFD defines correct detection based on the OSPA distance, which may occasionally obtain the output track where false alarms are present and the number of target observations is less than $M$. According to the definition of $P_{D,UPB}$, this does not count as a correct detection. However, the compliance of the condition shown in (27) indicates that the false alarms in the track do not affect subsequent target tracking, leading the consideration of correct detection. In practical engineering applications, such high-quality false alarms can also be advantageous for target track continuity.
@@
To completely show the experimental scenario and the performance of GLP-MFD, the processing results for one exemplar run are shown in Fig. 9. The observation area of interest is from $100$ km to $300$ km and from $-60^{\circ}$ to $60^{\circ}$ containing a total of 16 targets. The detection result corresponding to the primary threshold is shown in Fig. 9(a), with red solid lines indicating the tracks with correct target states. The processing results of the proposed GLP-MFD is displayed in Fig. 9(b), where the red solid lines represent the target tracks output from the network, black solid points indicate the false tracks, and blue squares denote the position measurements of targets.
@@
In the processing result, 3 false tracks and 14 target tracks are detected, while two target tracks are not detected. The retention of false tracks may result from their alignment with the multi-frame correlation characteristics of targets. However, they can be further eliminated through false track discrimination methods. Among 14 detected target tracks, 13 are identified as correct target tracks. The enlarged track in the figure has a large OSPA distance from the true target track due to the presence of false alarms, which affects the target track direction and may impact subsequent tracking. Therefore, it cannot be considered correctly detected. Nevertheless, these erroneous target tracks do not impose excessive negative effects. Over continuous multi-frame processing, the incorrect detection will gradually be filtered out, while the correct target observations within the erroneous tracks will remain and appear in the newly detected correct tracks. In summary, the proposed GLP-MFD algorithm demonstrates better detection performance for weak targets and has the ability to accurately detect target tracks while effectively suppressing false tracks.
5.4 Interpretability analysis
@@
To explore how MFLPN performs integrated track search and track detection based on input features, we analyze the importance of each input feature on the network output. The input features include signal-to-noise ratio (SNR), Doppler channel (DC), chronological information (CI), range-Doppler map (RDM), spatio-temporal coupling (STC) and the differences in Doppler channels (DCD). To analyze the contribution of these features to the network, techniques such as feature occlusion and feature permutation can be employed to eliminate the dependency between samples and their corresponding features [59],[60],[61]. The decline in network performance can be observed to confirm the importance of each feature. The core task of this work is the link prediction task, which determines whether the associations between observations originate from targets. To clarify the network decision criteria under different conditions, we use the extent of accuracy decline in link predictions caused by eliminating sample-feature dependencies as a reference metric. The box plots are used to explain the feature importance for the network [62]. The results of association decision in high SNR targets, low SNR targets, and false alarms are shown in Fig. 10.
@@
The results indicate that for high SNR targets, SNR is the most significant feature influencing performance. This is because a higher SNR enhances the separability between targets and false alarms, making it a key characteristic for distinguishing them. The next important features are the spatio-temporal coupling and the differences in Doppler channels. Effectively utilizing these features can enhance the accuracy of target detection. For low SNR targets, although SNR remains the most influential feature, its impact reduces compared to high SNR targets. To compensate, the network increasingly relies on the range-Doppler map, which reflects not only the intensity of echo signals, but also the distribution of target echo energy across the range and Doppler dimensions. This signal structure information facilitates deeper feature extraction by the network. Furthermore, since the network accuracy in making association decisions for low SNR targets is lower than that for high SNR targets, the range of outliers tends to be larger. In the case of associations between false alarms, the information extracted from SNR of noise measurements is limited. Therefore, the network relies more on the range-Doppler map, which contains rich signal structure information. Additionally, due to the randomness in the distribution of false alarms, the contribution of the differences in Doppler channels decreases, while the spatio-temporal coupling becomes critical for the network decision-making. Specifically, if the deviation in the coupling between measured Doppler and position information is small, the association is more likely to stem from a target. Conversely, it is more likely to originate from false alarms.
@@
Thus, through extensive data training, the proposed MFLPN can automatically adjust its association criteria based on data features under different conditions, thereby adaptively selecting corresponding features to make association decisions for observations from various sources. Compared to traditional model-driven algorithms, the proposed algorithm effectively integrates multi-dimensional information on different attributes to achieve unified track search and track detection in multiple frames.
@@ 6 SUMMARY
In this work, we proposed the GLP-MFD algorithm to cope with the weak target detection problem. By constructing observation association graphs with the maximum velocity constraint, the proposed algorithm ensured the successful target associations to mitigate the performance loss due to association failure. The MFLPN was designed based on graph neural networks to perform the link prediction task in graphs. The network fully integrated multi-dimensional information such as Doppler, signal structures, and spatio-temporal coupling, leading to accurate association decisions. In addition, the proposed algorithm completed track search and track detection in a unified manner, simplifying the processing logic while directly outputting target tracks. Numerical results indicated that the proposed algorithm enhanced target detection performance while suppressing false tracks, outperforming traditional single-frame and multi-frame detection methods. Moreover, the interpretability analysis revealed that the designed network effectively integrated the input features, enabling precise differentiation between target associations and false associations.
To further validate the robustness and versatility of the proposed algorithm, future work will explore its application in more complex environments, including those with interference, clutter, and varying noise levels. Additionally, although the computational efficiency of the designed network can be improved through GPU parallel processing, in practical scenarios involving a very large number of detection cells, the graph scale will become substantial. In such cases, how to reduce the computational complexity of the proposed algorithm requires further exploration.